The formula y=0.4x2-36x+1000 models the number of accidents, y, per 100 millions miles driven, for drivers x years old, where 20 < x< 50.Find the coordinates of the graph’s vertex and describe what this represent in practical terms.
· A company uses the function C(x)=0.01x2-1.2x+1000 to model the unit cost in dollars for producing x motors. Answer the following question; 1: For what number of motors is the unit cost at its minimum? 2: What happens if you produce 1 motor more or 1 motor less what the number of motors found in the first part of the problem. Discuss the effect of this change with your learning team.


Answer by stanbon(57967)

The formula y=0.4x^2-36x+1000 models the number of accidents, y, per 100 millions miles driven, for drivers x years old, where 20 < x< 50.
-------------------
Find the coordinates of the graph’s vertex and describe what this represent in practical terms.
---
The vertex occurs where x = -b/(2a) = 36/(2*0.4) = 45
and f(45) = 0.4(45)^2-36(45)+1000 = 190
Vertex:(45,190) means 190 accidents per 100 million miles of driving
for drivers 45 years old.
================================================

A company uses the function C(x)=0.01x^2-1.2x+1000 to model the unit cost in dollars for producing x motors. Answer the following question;
--------
1: For what number of motors is the unit cost at its minimum?
Minimum when x = -b/(2a) = 1.2/(2(0.01)) = 1.2/0.02 = 60
C(60) = 0.01(60)^2-1.2(60)+1000 = 964
Ans: # of motors = 60
-----------------
2: What happens if you produce 1 motor more or 1 motor less what the number of motors found in the first part of the problem.
--
Ans: Find C(59) = 964.01 and C(61) = 964.01
=====================================
Cheers,
Stan H.
==============